# Deep Time Series Models: A Comprehensive Survey and Benchmark

### v1

@@ 3.2.3 Matrix Factorization
The above-mentioned two decomposition methods are proposed for univariate series or applied to multivariate series in a variate-independent way. Here, we discuss a factorization-based decomposition for multivariate series. Specifically, many multivariate time series data in real-world scenarios can also be referred to as high-dimensional data. They can be formalized in the form of a matrix, whose rows correspond to variate and columns correspond to time points. Since variables in multivariate time series tend to be highly correlated, it can be possibly reduced to a more compact space. Matrix factorization methods [60] work by decomposing the high-dimensional series data into the product of two matrices in a lower-dimensional latent space. For a multivariate time series , as shown in Figure 2 the matrix can be approximated by the multiplications of two lower rank embedding matrix, , in which , and is a hyperparameter.
Besides the estimation, there are regularizers to avoid overfitting problems in factorization. Going beyond the canonical design that takes the squared Frobenius norm as regularizers, Temporal regularized matrix factorization (TRMF) [61] designs an autoregressive-based temporal regularizer to describe temporal dependencies among latent temporal embeddings. Further, [62] extended TRMF with a new spatial autoregressive regularizer to estimate low-rank latent factors by simultaneously learning the spatial and temporal autocorrelations. NoTMF [63] integrates the vector autoregressive process with differencing operations into the classical low-rank matrix factorization framework to better model real-world time series data with trend and seasonality. Eliminating the need for tuning regularization parameters, BTF [64] is a fully Bayesian model that integrates the probabilistic matrix factorization and vector autoregressive process into a single probabilistic graphical model. Instead of using an autoregressive-based temporal regularization, DeepGLO [65] utilizes a temporal convolution network for regularization to capture non-linear dependencies. LSTM-GL-ReMF [66] contains an LSTM-based temporal regularizer to learn complex long-term and short-term non-linear temporal correlations and a Graph Laplacian spatial regularizer [67] to capture spatial correlations.
@@ 3.3 Fourier Analysis
Fourier analysis [68, 69] can convert a physical signal into the Fourier domain to highlight the inherent frequency properties of the original data and has been a well-acknowledged analysis tool in extensive areas. Since time series are usually recorded as a sequence of discrete time points by sampling the original continuous signals, Fourier analysis has become one of the mainstream tools in time series modeling and has been demonstrated favorable effectiveness and efficiency [70, 71]. Introducing the Fourier domain not only augments the representation of the original series but also provides a global view since the frequency spectrum distribution, which can indicate essential periodic properties of time series. In practice, Fast Fourier Transform (FFT) [72] and Wavelet Transform (WT) [73] as the basic algorithms connecting the discrete temporal domain to the frequency domain, have gained increasing popularity in the modular design of deep time series models [74, 75, 76, 77, 78, 79, 80, 81]. Existing approaches can be roughly divided into two categories: time-domain and frequency-domain modeling.
@@ 3.3.1 Time-Domain Modeling
The fundamental principle behind the Fourier transform is that sequential data can be decomposed and represented by a series of periodic signals. Consequently, it can be used to identify potentially dominant periods and their corresponding frequencies in the data by analyzing the highest amplitude components. As a typical practice, TimesNet [3] employs the Fast Fourier Transform (FFT) to extract the most significant frequencies with the highest amplitude values, subsequently reshaping the 1D time series data into a 2D space based on the identified periods for better representation learning. Following TimesNet, PDF [82] posits that frequencies with larger values facilitate a more discernible distinction between long-term and short-term relationships.
In addition to exploiting the information of the sequence obtained by the Fourier Transformer, some works attempt to perform efficient computation through the Fast Fourier Transformer. Auto-correlation is a fundamental concept in time series analysis that measures the dependence between observations at different time points within a sequence of data. The Wiener-Khinchin theorem [83] provides a mathematical relationship between the auto-correlation function and the power spectral density (PSD) of a stationary random process, where the auto-correlation function represents the inverse Fourier transform of the PSD. Taking the data as a real discrete-time process, Autoformer [22] proposes an Auto-Correlation mechanism with an efficient Fast Fourier Transforms to capture the series-wise correlation.
The frequency-domain representation provides information about the amplitudes and phases, where low-frequency components correspond to slower variations or trends in the signal, and high-frequency components capture fine details or rapid variations. A significant body of work has focused on leveraging frequency-domain information to enhance the model’s capability in capturing temporal dependencies. FiLM [84] introduces Frequency Enhanced Layers (FEL) which combine Fourier analysis with low-rank approximation to keep the part of the representation related to low-frequency Fourier components and the top eigenspace to effectively reduce the noise and boost the training speed. FITS [85] integrates a low-pass filter (LPF) to eliminate high-frequency components above a specified cutoff frequency, thereby compressing the model size while preserving essential information. From an opposite idea, FEDformer [86] posits that retaining only low-frequency components is insufficient for time series modeling, as it may dismiss important fluctuations in the data. Based on the above considerations, to capture the global view of time series, FEDformer represents the series by randomly selecting a constant number of Fourier components, including both high-frequency and low-frequency components.
@@ 3.3.2 Frequency-Domain Modeling
Building on time-frequency analysis in signal processing, several approaches have been developed to study time series simultaneously in both the time and frequency domains. ATFN [87] comprises an augmented sequence-to-sequence model that learns the trending features of complex non-stationary time series, along with a frequency-domain block designed to capture dynamic and intricate periodic patterns. TFAD [88] introduces a time-frequency analysis-based model that employs temporal convolutional networks to learn both time-domain and frequency-domain representations.
Some works have developed specialized deep learning architecture to process the frequency domain of time series. STFNet [75] applies Short-Time Fourier Transform to input signals and applies filtering, convolution, and pooling operations directly in the frequency domain. StemGNN [28] combines Graph Fourier Transform (GFT) and Discrete Fourier Transform to model both inter-series correlations and temporal dependencies. EV-FGN [89] uses a 2D discrete Fourier transform on the spatial-temporal plane of the embeddings and performs graph convolutions for capturing the spatial-temporal dependencies simultaneously in the frequency domain. FreTS [90] leverages Discrete Fourier Transform (DFT) to transform the data into the frequency domain spectrum and introduces frequency domain MLPs designed for complex numbers with separated modeling for the real parts and the imaginary parts. FCVAE [91] integrates both the global and local frequency features into the condition of Conditional Variational Autoencoder (CVAE) concurrently. Recent TSLANet [92] propose a lightweight Adaptive Spectral Block (ASB) to replace the self-attention mechanism, which is achieved via Fourier-based multiplications by global and local filters. FourierDiffusion [93] explores extending the score-based SDE formulation of diffusion to complex-valued data and therefore implements time series diffusion in the frequency domain.
@@ 4 Model Architectures
As we have discussed in Section 2, the time series model needs to unearth the intrinsic temporal dependencies and variate correlations that lie in observations. In this section, we provide a technical review of the existing deep time series models. As we have presented in Figure 3, existing works can be classified into five categories based on their backbone architecture, namely MLP-based, RNN-based, CNN-based, GNN-based, and Transformer-based.
4.1 Multi-Layer Perceptrons
@@
As a representation of traditional statistical time series models, the Auto-regressive (AR) model assumes that the model output depends linearly on its own historical values. Inspired by the remarkable performance of auto-regressive models, Multi-Layer Perceptrons (MLP) have become a popular architecture for modeling time series data.
@@
As a representative work of linear-based models, N-BEATS [24] is a pure MLP-based deep time series model without any time-series-specific knowledge to capture the temporal patterns in time series. Specifically, as described in Equ. (4) N-BEATS consists of deep stacks of fully-connected layers with two residual branches in each layer, one is for the backcast prediction and the other one is the forecast branch. Extending the idea of neural basis expansion analysis, N-HiTs [57] use multi-rate signal sampling and hierarchical interpolation and N-BEATSx [94] incorporate exogenous variables to enhance the prediction.
@@
Recent research by DLinear [52], also referred to as LTSF-Linear, challenges the effectiveness of complicated deep architecture in temporal modeling. It argues a simple linear regression in the raw space that achieves remarkable performance in both modeling and efficiency. As illustrated in Figure 4, prevalent MLP-based deep time series models consist of simple linear layers primarily designed for forecasting tasks. Also lightweight but effective, FITS [85] advocates time series analysis can be treated as interpolation exercises within the complex frequency domain and further introduces a complex-valued linear layer to learn amplitude scaling and phase shift in the frequency domain. Inspired by MLP-Mixer [95] in computer vision, several works have attempted to utilize MLPs to model both temporal and variate dependencies. TSMixer [96] contains interleaving time-mixing and feature-mixing MLPs to extract information from different perspectives. To better model the global dependencies in time series data, FreTS [90] investigates the learned patterns of frequency-domain MLPs which are operated on both inter-series and intra-series scales to capture channel-wise and time-wise dependencies in multivariate data.
@@
Recent works have moved beyond using simple linear layers over discrete time points. TimeMixer suggests that time series exhibit distinct patterns in different sampling scales and proposes an MLP-based multiscale mixing architecture. TiDE [97] incorporates exogenous variables to enhance the time series prediction. Based on Koopman theory and Dynamic Mode Decomposition (DMD) [98], which is a dominant approach for analyzing complicated dynamical systems, Koopa [99] hierarchically disentangles dynamics through an end-to-end predictive training framework and can utilize real-time incoming series for online development.
@@ 4.2 Recurrent Neural Networks
Recurrent Neural Networks (RNNs) are specifically designed to model sequential data [100, 101, 102], such as natural language processing [103] and audio modeling [104]. Since time series are also serial in nature, RNNs have emerged as a popular choice for analyzing time series data [105]. Existing RNN-based deep time series models focus on combating the gradient vanishing problem caused by the vanilla recurrent structure and modeling the mutual correlation among multivariate variables. Previous works [106, 107, 108, 109, 110] use variants of RNN to model temporal dependencies.
@@
LSTNet [111] combines the recurrent structure with the convolutional layer to capture both the short-term local dependency between variables and long-term patterns for time series. Moreover, a novel recurrent-skip component based on the periodic pattern is introduced to alleviate gradient vanishing in modeling long-term dependencies. Similarly, DA-RNN [112] combines the recurrent unit with a dual-stage attention mechanism to adaptively extract relevant series at each time step. Beyond deterministic forecasts, DeepAR [113] proposes an auto-regressive recurrent network model to predict the probability distribution of further time points. Technologically, it learns not only the seasonal behavior with time series but dependencies on given covariates across time series, allowing the model to make predictions even when there is little or no historical data.
Also based on Markovian state representation, the State Space Model (SSM) [114] is another classical mathematical framework that captures the probabilistic dependence between observed measurements in stochastic dynamical systems. Concretely, a single-input single-output (SISO) linear state space model is defined as follows:
where are input signal, latent state, and output signal respectively. The system is characterized by the matrices can be learned by the deep neural network.
@@
SSMs have proven their effectiveness and efficiency in processing well-structured time series data, but traditional approaches have to refit each time series sample separately and therefore cannot infer shared patterns from a dataset of similar time series. With the rise of deep learning models, modern SSMs are often implemented in a recurrent manner. By adapting and propagating a deterministic hidden state, RNNs are able to represent long-term dependencies in continuous data which offer an alternative to classical state space models. Therefore, some work [115, 116] have attempted to fuse classical state space models with deep neural networks. Representative like Deep State Spaces Model (DSSM) [117], using a recurrent neural network (RNN) to parametrize a particular linear SSM, takes advantage of incorporating structural assumptions and learning complex patterns. Structured State Space sequence model (S4) [118] introduces a new parameterization for the SSM by conditioning matrix with a low-rank correction, allowing it to be diagonalized stably, which empowers the model with better long-term modeling capacity. Similar to S4, LS4 [119] is a generative model with latent space evolution following a state space ordinary differential equations (ODE).
Recent work on Mamba [120] has emerged as a powerful method for modeling long-context sequential data while scaling linearly with sequence length. Utilizing a simple selection mechanism that parameterizes the SSM parameters based on the input, Mamba can discern the importance of information in a manner similar to the attention mechanism, posing a potentially effective way to sequential modeling.
@@ 4.3 Convolutional Neural Networks
Since the semantic information of time series is mainly hidden in the temporal variation, Convolutional neural networks (CNN) [18, 121] have become a competitive backbone for their ability to capture local features and pattern recognition. By leveraging convolutions and hierarchical feature extraction, CNNs have shown remarkable success in various computer vision tasks, such as image classification [122], segmentation [123] and object detection [124].
Considering the temporal continuity of time series data, previous works [125, 126, 127] apply one-dimensional CNN (1D CNN) to capture the local patterns of time series data. Recent SCINet[128] applies normal convolutions with a hierarchical downsample-convolve-interact architecture to capture dynamic temporal dependencies at different temporal resolutions of time series data. Inspired by the idea of masked convolution [129], Wavenet[130] introduces causal convolution and dilated causal convolution to model long-range temporal causality. Similar to Wavenet, Temporal Convolutional Networks (TCN) [131] uses a stack of dilated convolutional kernels with progressively enlarged dilation factors to achieve a large receptive field. However, the limited receptive field of TCN makes it difficult for them to capture global relationships in time series data. Based on TCN, MICN[132] is a local-global convolution network that combines different convolution kernels to model temporal correlation from a local and global perspective. ModernTCN [133] boosts the traditional TCN to capture cross-time and cross-variable dependency by DWConv and ConvFFN separately. Considering that DWConv is proposed to learn temporal information, it is operated variate-independently to learn the temporal dependency of each univariate time series.
Beyond 1D space, motivated by the periodicity properties of time series data, TimesNet [3] transforms the 1D time series data into a set of 2D tensors in each TimesBlock based on the estimated period lengths, where the inter-period variations are presented in tensor columns and inner-period ones are shown in tensor rows. Here is a hyperparameter, corresponding to multiple 1D-to-2D transformations with different periods. Then it applies inception block [134, 135] to process the transformed 2D tensors, which can be summarized as:

### v2

@@ 3.2.3 Matrix Factorization
The above-mentioned two decomposition methods primarily address univariate series or are applied to multivariate series in a variate-independent way. Here, we explore factorization-based decomposition that is particularly pertinent to multivariate time series. These high-dimensional data can be formalized as a matrix where rows represent variates and columns denote time points. Since variables in multivariate time series tend to be highly correlated, they can be reduced to a more compact space. Matrix factorization methods [58] achieve this by decomposing the high-dimensional series data into the product of two matrices in a lower-dimensional latent space. For a multivariate time series , the matrix can be approximated by the multiplications of two lower rank embedding matrix, , in which , and is a hyperparameter.
Besides the estimation, regularizers are critical in factorization-based models to mitigate overfitting. Going beyond the canonical design that takes the squared Frobenius norm as regularizers, temporal regularized matrix factorization (TRMF) [59] designs an autoregressive-based temporal regularizer to describe temporal dependencies among latent temporal embeddings. Further, [60] extended TRMF with a new spatial autoregressive regularizer to estimate low-rank latent factors by simultaneously learning the spatial and temporal autocorrelations. For enhanced modeling of the trends and seasonality of real-world time series data, NoTMF [61] integrates the vector autoregressive process with differencing operations into the classical low-rank matrix factorization framework. Addressing the challenge of parameter tuning, BTF [62] offers a fully Bayesian approach, uniting probabilistic matrix factorization and the vector autoregressive process within a single probabilistic graphical model. Diverging from autoregressive-based temporal regularization, DeepGLO [63] leverages a temporal convolution network to capture non-linear dependencies. Similarly, LSTM-GL-ReMF [64] employs an LSTM-based temporal regularizer for complex long-term and short-term non-linear temporal correlations, complemented by a graph Laplacian spatial regularizer [65] for capturing spatial correlations.
@@ 3.3 Fourier Analysis
Fourier analysis [66, 67] is a powerful mathematical technique that transforms a physical signal into the frequency domain, thereby illuminating the inherent periodic properties of the original data. It has long been recognized as a fundamental analytical tool across diverse domains. Given that time series data are typically recorded as sequences of discrete observations sampled from continuous signals, Fourier analysis has naturally emerged as a mainstream approach in time series modeling with favorable effectiveness and efficiency [68, 69]. Introducing the Fourier domain not only enhances the representation of the original series but also provides a global perspective on temporal dynamics. The resulting frequency spectrum distribution offers crucial insights into the essential periodic components and underlying cyclical behaviors of time series. In practice, algorithms such as Fast Fourier Transform (FFT) [70] and Wavelet Transform (WT) [71] act as foundational bridges, connecting the discrete temporal domain to the frequency domain. These techniques have gained increasing popularity in the modular design of deep time series models [72, 73, 74, 75, 76, 77, 78, 79]. Existing approaches can be broadly categorized into two main types: time-domain and frequency-domain modeling.
@@ 3.3.1 Time-Domain Modeling
The fundamental principle behind the Fourier transform is that sequential data can be decomposed and represented by a series of periodic signals. Consequently, it can be used to identify potentially dominant periods and their corresponding frequencies in the data by analyzing the highest amplitude components. As a prominent integration, TimesNet [3] employs the Fast Fourier Transform (FFT) to extract the most significant frequencies with the highest amplitude values. It then reshapes the 1D time series data into a 2D space based on these identified periods, thereby facilitating hierarchical representation learning. Building upon TimesNet, PDF [80] posits that frequencies with larger values facilitate a more discernible distinction between long-term and short-term relationships.
The frequency-domain representation provides valuable information regarding both amplitudes and phases. Low-frequency components, for instance, correspond to slower variations or underlying trends in the signal, while high-frequency components capture finer details or rapid fluctuations. A significant body of work has focused on leveraging frequency-domain information to enhance the model’s capability in capturing temporal dependencies. FiLM [81] combines Fourier analysis with low-rank approximation and introduces a frequency enhanced layer (FEL). This approach retains portions of the representation related to low-frequency Fourier components and the top eigenspace, thereby effectively reducing noise and accelerating training speed. FITS [82] integrates a low-pass filter (LPF) to eliminate high-frequency components above a specified cutoff frequency, thereby compressing the model size while preserving essential information. Taking an opposing view, FEDformer [83] posits that retaining only low-frequency components is insufficient for time series modeling, as it risks dismissing important data fluctuations. Based on this consideration, FEDformer aims to capture a global view of time series by representing the series using a randomly selected constant number of Fourier components, encompassing both high-frequency and low-frequency contributions.
Beyond merely exploiting sequence information obtained from Fourier transforms, some research focuses on achieving efficient computation through the Fast Fourier Transform. Auto-correlation is a fundamental concept in time series analysis that measures the dependence between observations at different time points within a sequence of data. The Wiener-Khinchin theorem [84] provides a mathematical relationship between the auto-correlation function and the power spectral density (PSD) of a stationary random process, where the auto-correlation function represents the inverse Fourier transform of the PSD. Taking the data as a real discrete-time process, Autoformer [22] proposes an Auto-Correlation mechanism with an efficient Fast Fourier Transform to capture the series-wise correlation.
@@ 3.3.2 Frequency-Domain Modeling
Building on advancements in time-frequency analysis from signal processing, several approaches have been developed to analyze time series simultaneously in both the time and frequency domains. ATFN [85], for instance, utilizes an augmented sequence-to-sequence model to learn trending features of non-stationary time series, alongside a frequency-domain block for capturing periodic patterns. TFAD [86] introduces a time-frequency analysis-based model that employs temporal convolutional networks to obtain both time-domain and frequency-domain representations.
Some works have developed specialized deep learning architecture to process the frequency domain of time series. STFNet [73] applies Short-Time Fourier Transform to input signals and applies filtering, convolution, and pooling operations directly in the frequency domain. StemGNN [87] combines Graph Fourier Transform (GFT) and Discrete Fourier Transform to model both inter-series correlations and temporal dependencies. EV-FGN [88] uses a 2D discrete Fourier transform on the spatial-temporal plane of the embeddings and performs graph convolutions for capturing the spatial-temporal dependencies simultaneously in the frequency domain. FreTS [89] leverages Discrete Fourier Transform (DFT) to transform the data into the frequency domain spectrum and introduces frequency domain MLPs designed for complex numbers with separated modeling for the real parts and the imaginary parts. FCVAE [90] integrates both the global and local frequency features into the condition of Conditional Variational Autoencoder (CVAE) concurrently. Recent TSLANet [91] proposes a lightweight Adaptive Spectral Block (ASB) to replace the self-attention mechanism, which is achieved via Fourier-based multiplications by global and local filters. FourierDiffusion [92] extends the score-based SDE formulation of diffusion models to complex-valued data and therefore implements time series diffusion in the frequency domain.
@@ 4 Model Architectures
As discussed in Section 2, time series models aim to unearth the intrinsic temporal dependencies and variate correlations to uncover underlying patterns and dynamics within observations. Existing approaches can be broadly categorized into five groups based on their architectural backbones (Figure 2), namely, MLP-based, RNN-based, CNN-based, GNN-based, and Transformer-based. Each of these architectures possesses distinct strengths and design principles that address specific challenges in time series analysis. In this section, we provide a comprehensive technical review of current deep time series models, specifically highlighting their distinct approaches to time series modeling.
4.1 Multilayer Perceptrons
@@
As a cornerstone of traditional statistical time series models, the Auto-regressive (AR) model posits that the model output depends linearly on its own past values. Inspired by the notable efficacy of auto-regressive models, Multilayer Perceptrons (MLP) have become a popular architecture for time series modeling.
@@
As a prime example of MLP-based models, N-BEATS [24] integrates no time-series-specific components to capture temporal patterns. Specifically, as described in Eq. (3), N-BEATS consists of deep stacks of fully-connected layers with two residual branches in each layer, one for the backcast prediction and the other for the forecast branch. Expanding the idea of neural basis expansion, N-HiTs [55] enhances prediction through multi-rate signal sampling and hierarchical interpolation, while N-BEATSx [93] further extends the framework by incorporating exogenous variables.
@@
Recent research, notably DLinear [50], rethinks the effectiveness of complicated deep architectures for temporal modeling. It argues a simple linear regression in the raw space that achieves remarkable performance in both modeling and efficiency. As illustrated in Figure 3, many prevalent MLP-based deep time series models, primarily designed for forecasting, rely on such simple linear layers. Also lightweight but effective, FITS [82] advocates that time series forecasting can be treated as interpolation exercises within the complex frequency domain. It introduces a complex-valued linear layer to learn amplitude scaling and phase shift directly in the frequency domain. Inspired by MLP-Mixer [94] in computer vision, several works have attempted to leverage MLPs to model both temporal and variate dependencies. TSMixer [95] contains interleaving time-mixing and feature-mixing MLPs to extract information from different perspectives. To better capture global dependencies, FreTS [89] investigates the learned patterns of frequency-domain MLPs, which operate on both inter-series and intra-series scales to capture channel-wise and time-wise dependencies in multivariate data.
@@
Moving beyond simple linear layers over discrete time points, TimeMixer suggests that time series exhibit distinct patterns across different sampling scales and introduces an MLP-based multiscale mixing architecture. TiDE [96] incorporates exogenous variables to enhance time series prediction. Based on Koopman theory and Dynamic Mode Decomposition (DMD) [97], which is a dominant approach for analyzing complicated dynamical systems, Koopa [98] hierarchically disentangles dynamics through an end-to-end predictive training framework and can utilize real-time incoming series for online development.
@@ 4.2 Recurrent Neural Networks
Recurrent Neural Networks (RNNs) are specifically designed for sequential data [99, 100, 101], such as natural language processing [102] and audio modeling [103]. Given the inherent sequential nature of time series, RNNs have naturally become a popular choice for time series analysis [104]. Existing RNN-based deep time series models primarily focus on combating the gradient vanishing problem caused by the vanilla recurrent structure and modeling the mutual correlation among multivariate variables.
@@
Numerous prior works [105, 106, 107, 108, 109] have leveraged various RNN variants to capture complex dependencies. LSTNet [110] combines the recurrent structure with the convolutional layer to capture both the short-term local dependency between variables and long-term patterns for time series. It further introduces a novel recurrent-skip component, based on periodic patterns, to alleviate gradient vanishing in modeling long-term dependencies. Similarly, DA-RNN [111] combines the recurrent unit with a dual-stage attention mechanism to adaptively extract relevant series at each time step. Beyond deterministic forecasts, DeepAR [112] proposes an auto-regressive recurrent network model to predict the probability distribution of further time points. Technologically, it learns not only the seasonal behavior with time series but dependencies on given covariates across time series, allowing the model to make predictions even when there is little or no historical data. Recent work [77] introduces a novel DTWAtt module, which integrates a data-dependent wavelet function into the BiLSTM network as wavelet attention to analyze dynamic frequency components in nonstationary time series.
@@
Based on Markovian state representation, the State Space Model (SSM) [113] is a classical mathematical framework for capturing probabilistic dependencies between observed measurements in stochastic dynamical systems. While SSMs have proven their effectiveness and efficiency in processing well-structured time series data, traditional approaches have to refit each series separately and therefore cannot infer shared patterns from a dataset of similar time series. With the advent of deep learning, modern SSMs are often implemented in a recurrent manner. By adapting and propagating a deterministic hidden state, RNNs offer an alternative to classical state space models by representing long-term dependencies in continuous data. Consequently, some works [114, 115] have sought to fuse classical state space models with deep neural networks. Representative among these, the Deep State Space Model (DSSM) [116] leverages a recurrent neural network (RNN) to parametrize a specific linear SSM, thereby incorporating structural assumptions while learning complex patterns. The Structured State Space sequence model (S4) [117] introduces a novel parametrization for the SSM by conditioning matrix with a low-rank correction, which allows for stable diagonalization and significantly enhances the model’s long-term modeling capacity. More recently, Mamba [118] has emerged as a powerful method for modeling long sequential data with linear complexity w.r.t. sequence length. Utilizing a simple selection mechanism that parameterizes the SSM parameters based on input, Mamba can discern valuable information in observations, posing an effective way for time series analysis.
Going beyond learning among discretized time steps, Neural Ordinary Differential Equations (Neural ODEs) [119] extend RNNs into a continuous-time version. Technically, Neural ODEs parameterize the derivative of hidden states through neural networks, enabling seamless modeling of temporal evolution in a differentiable manner. Given the ability to model continuous-time, several works have explored applying Neural ODEs to handle irregularly-sampled time series data [120, 121]. Building on these advancements, [122] integrates Neural ODEs with physical priors to augment physical models with deep networks for forecasting complex dynamics. These capabilities make Neural ODEs a compelling complement to RNN-based methods, offering a more flexible and robust approach for time series analysis.
@@ 4.3 Convolutional Neural Networks
Since the semantic information of time series is mainly hidden in the temporal variation, Convolutional neural networks (CNN) [18, 123] have become a competitive backbone. Their strength lies in their ability to capture local features and recognize intricate patterns, which has been demonstrated with remarkable success in various computer vision tasks, including image classification [124], segmentation [125], and object detection [126].
Considering the temporal continuity of time series data, previous works [127, 128, 129] apply a one-dimensional CNN (1D CNN) to discern the local patterns of time series data. Recent SCINet[130] employs canonical convolutions with a hierarchical downsample-convolve-interact architecture to capture dynamic temporal dependencies across varying temporal resolutions. Inspired by the idea of masked convolution [25], Wavenet[131] introduces causal convolution and dilated causal convolution to model long-range temporal causality. Similarly, Temporal Convolutional Networks (TCN) [132] uses a stack of dilated convolutional kernels with progressively enlarged dilation factors to achieve a large receptive field. However, the limited receptive field of TCN often hinders its capacity to capture global relationships in time series data. Based on TCN, MICN[133] is a local-global convolution network that combines different convolution kernels to model temporal correlation from a local and global perspective. ModernTCN [134] enhances the traditional TCN by separately employing Depthwise Separable Convolutions (DWConv) and Convolutional Feed-Forward Networks (ConvFFN) to capture cross-time and cross-variable dependencies, respectively.
Motivated by the inherent periodicity of time series data, TimesNet [3] goes beyond traditional 1D modeling. It [3] transforms the 1D time series data into a set of 2D tensors in each TimesBlock based on the estimated period lengths, where the inter-period variations are presented in tensor columns and inner-period ones are shown in tensor rows. Here is a hyperparameter, corresponding to multiple 1D-to-2D transformations with different periods. Subsequently, TimesNet applies an Inception Block [135, 136] to process these transformed 2D tensors, which can be summarized as: